OpenStudy (anonymous):

4(6+x) is greater than or equal to -2x

OpenStudy (anonymous):

solve for x?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

x greater than or equal to -4

OpenStudy (anonymous):

Remember to get all xs on one side, then everything else on the other.

OpenStudy (mathteacher1729):

\[4(6+x) \geq -2x\] \[24 + 4x \geq -2x\] \[24 +6x \geq 0\] \[6x \geq -24\] \[x \geq -4\] So any value of \(x\) which is greater than or equal to \(-4\) should make the original inequality true. Let's try it out with x = -4 4(6 + (-4)) = 4 ( 2) = 8 and -2(-4) = -8 It is indeed true that 8 is greater than or equal to -8. Let's try it out for x = 0 4(6+(0)) = 24 and -2(0) = 0 and one again 24 is greater than or equal to 0.